OpenStudy (anonymous):

find the value expression 12-(-13)+(-1)

OpenStudy (anonymous):

it would change to 12+13+(-1) which is really just 12+13-1....12+13=25-1=24

OpenStudy (saifoo.khan):

hard as diamond.

OpenStudy (anonymous):

then answer is 26 because 12-(-13)=25+(-1)=26

OpenStudy (anonymous):

-1 actually mean to subtract 1 so you subrtact 1 from 25 which is 24

OpenStudy (anonymous):

subtract*

OpenStudy (saifoo.khan):

Meera's right!

OpenStudy (anonymous):

do it in a cal. and the answer is 26

OpenStudy (saifoo.khan):

25+(-1) 25 - 1 24

OpenStudy (anonymous):

its 24 sweetheart =]

OpenStudy (saifoo.khan):

yes it is.

OpenStudy (anonymous):

ok